# QServe: W4A8KV4 Quantization and System Co-design for Efficient LLM Serving

### v2

@@ IV-B SmoothAttention
As illustrated in Figure 16, directly reducing the KV cache to 4 bits significantly degrades the LLM accuracy. We visualize the magnitude distributions of the sampled Key and Value cache activations in Figure 7. We observe that: the Value matrices show no significant outlier pattern, whereas Key matrices tend to have fixed outlier channels in each head. These outliers are $\sim$10$\times$ larger than most of activation values. Though they can be easily handled KV8 quantization in prior works [38], it places challenging obstacle to KV4 quantization due to less quantization levels.
Inspired by SmoothQuant [38], we propose SmoothAttention to scale down the outlier channels in Key cache by a per-channel factor $\mathbf{\lambda}$:
$\small\mathbf{Z}=\left(\mathbf{Q}\mathbf{\Lambda}\right)\cdot\left(\mathbf{K}\mathbf{\Lambda}^{-1}\right)^{T},\;\;\;\mathbf{\Lambda}=\mathrm{diag}\left(\mathbf{\lambda}\right)$ (7)
@@
$\small\mathbf{\lambda}_{i}=\max\left(|\mathbf{K}_{i}|\right)^{\alpha}.$ (8)
@@
$\small\mathbf{\lambda}_{i}=\lambda_{i+\frac{D}{2}}=\max\left(\max\left(|\mathbf{K}_{i}|\right),\max\left(|\mathbf{K}_{i+\frac{D}{2}}|\right)\right)^{\alpha}$ (9)
@@
IV-C General LLM Quantization Optimizations
@@
IV-C1 Block Input Module Rotation
In transformer blocks, we define the components that take in the block inputs as input modules, such as the QKV Projection Layer and the FFN 1st Layer. As shown in Figure 9, inspired by [2, 4], we rotate the block input activations by multiplying the rotation matrix. To keep mathematical equivalence of linear layers, we rotate the corresponding weights accordingly in the reversed direction. After rotation, each channel’s activations are linear combinations of all other channels, and thus outlier channels are effectively suppressed. Furthermore, since rotation is a unitary transformation, we can fuse the rotation matrix with the previous linear layers’ weights. We simply choose the scaled Hadamard matrix as the rotation matrix.
IV-C2 Block Output Module Smoothing
Output modules refer to those layers that generate block outputs, such as the Output Projection Layer and FFN 2nd Layer. As shown in Figure 9, inspired by [38], we smooth the block intermediate activations through dividing them by a per-channel smoothing factor. Original SmoothQuant does not smooth the block intermediate activations; moreover, if we directly smooth these modules with the same migration strength as input modules (e.g., q_proj, up_proj), the evaluated Wikitext-2 perplexity of the Llama-2-7B model will drop by as much as 0.05. In practice, we find that the migration strength $\alpha$ should be near 0. That is, the smoothing factor $\lambda$ is mostly determined by weights instead of activations, which is very different from the observations in SmoothQuant.
IV-C3 Activation-Aware Channel Reordering
@@
Both AWQ [23] and Atom [44] have observed that maintaining the salient weights in FP16 can significantly improve model accuracy. These salient weights can be identified by the activation distribution. Instead of introducing mixed-precision quantization used by Atom, we propose activation-aware channel reordering as shown in Figure 10. We use $\max\left(|\mathbf{X}|\right)$ to determine the channel salience, and then reorder channels so that channels with similar salience are in the same quantization group.
IV-C4 Weight Clipping
Weight clipping is another popular quantization optimization technique. It applies a clip ratio $\alpha$ to the dynamic range in Equation 2 by letting $\mathbf{W}_{\max}=\alpha\max\left(\mathbf{W}\right)$ and $\mathbf{W}_{\min}=\alpha\min\left(\mathbf{W}\right)$. Previous approaches [12, 44, 2, 23] grid search the clip ratio $\alpha$ to minimize either quantization error of tensor itself (i.e., $\|\mathbf{W}-Q\left(\mathbf{W};\alpha\right)\|$) or output mean square error (i.e., $\|\mathbf{X}\mathbf{W}^{T}-\mathbf{X}Q\left(\mathbf{W}^{T};\alpha\right)\|$. In QServe, we minimize the layer output error for all linear layers, expect for q_proj and k_proj, for which we optimize block output mean square error:
$\small\arg\min_{\alpha}\|\mathrm{Block}\left(\mathbf{X};\mathbf{W}\right)-\mathrm{Block}\left(\mathbf{X};Q\left(\mathbf{W};\alpha\right)\right)\|.$ (10)
V QServe Serving System
To this end, we have presented the QoQ quantization algorithm, which aims to minimize accuracy loss incurred by W4A8KV4 quantization. However, realizing the theoretical throughput benefits in Figure 3 remains challenging. Thus, in this section, we will delve into the QServe system design, which is guided by two important principles: I. Reducing main loop overhead in GEMM kernels; II. Making fused attention kernels memory bound.
V-A QServe System Runtime
@@
KV Cache Management. To avoid memory fragmentation, we follow vLLM [21] and TensorRT-LLM [25] to adopt paged KV caches. In contrast to these two frameworks, which perform per-tensor, static quantization (i.e., scaling factors computed offline) on KV caches, QServe requires per-head, dynamic KV quantization to maintain competitive accuracy due to the lower bit precision (4 vs. 8). We therefore store FP16 scaling factors and zero points for each head immediately following the quantized KV features in each KV cache page, allowing these values to be updated on-the-fly. QServe also supports in-flight batching, similar to vLLM and TensorRT-LLM.
V-B W4A8 GEMM in QServe
As discussed in Section III, the main loop overhead poses a significant obstacle in allowing quantized GEMMs to attain the theoretical performance gains projected by the roofline model (Figure 3). Therefore, the focus of QServe W4A8 GEMM is to reduce main loop overhead. Specifically, we address the costs of pointer arithmetic operations through compute-aware weight reorder, and reduce dequantization overhead through a subtraction after multiplication computation order and register-level parallelism.
V-B1 Compute-Aware Weight Reorder
@@
V-B2 Fast Dequantization in Per-Channel W4A8 GEMM
@@
$\small\mathbf{O}=\hat{\mathbf{X}}\hat{\mathbf{W}}=(\mathbf{Q}_{\mathbf{X}}\odot\mathbf{S}_{\mathbf{X}})((\mathbf{Q}_{\mathbf{W}}-\mathbf{Z}_{\mathbf{W}})\odot\mathbf{S}_{\mathbf{W}}),$ (11)
@@
$\small\mathbf{O}$ $=(\mathbf{Q}_{\mathbf{X}}\odot\mathbf{S}_{\mathbf{X}})(\mathbf{Q}_{\mathbf{W}}\odot\mathbf{S}_{\mathbf{W}}-\mathbf{ZS}_{\mathbf{W}})$ (12)
@@
V-B3 Fast Dequantization in Per-Group W4A8 GEMM
@@ V-B4 General Optimizations
V-C KV4 Attention in QServe
TABLE I: A naive KV4 attention implementation is 1.7$\times$ faster on L40S than TRT-LLM-KV8, but is 1.1-1.2$\times$ slower on A100 due to earlier CUDA core roofline turning point.
@@ V-C KV4 Attention in QServe
We start with the KV8-attention decoding stage kernel from TensorRT-LLM as our baseline and replace all static, per-tensor quantized 8-bit KV cache accesses and conversions with their dynamic, per-head quantized 4-bit counterparts. This direct replacement immediately leads to 1.7$\times$ speedup on L40S, but results in 1.2$\times$ slowdown on A100 (Table I), compared to the KV8 baseline.
@@
Once again, our analysis reveals that the devil is in the slow CUDA cores, which are responsible for executing the attention kernels during the decoding stage. While each individual batched GEMV has a computation intensity of 1 MAC / element, the computation intensity escalates significantly for a fused attention kernel that combines all the arithmetics and KV cache updates. As an illustration, naively dequantizing a single INT4 number from the KV cache necessitates 5 ALU Ops. This includes mask and shift operations to isolate the operand, type conversion from integer to floating-point representation, and floating point mul and sub to obtain the final results. It is crucial to note that the roofline turning point for A100 FP32 CUDA cores is merely 9.8 Ops/Byte. That said, the dequantization of KV operands alone already saturates this bound, leading to the surprising observation that the fused KV4 attention kernel can become compute-bound on datacenter GPUs like A100. In fact, similar observations hold in other systems like QuaRot [2] and Atom [44]. Specifically, QuaRot introduces compute-intensive Hadamard transformation [4] in the attention operator, making it hard to achieve real speedup over TRT-LLM-KV8 with 4-bit quantized KV caches.
@@
To mitigate the compute-bound bottleneck, it is important to shift the decoding stage KV4 attention kernels away from the compute-bound region. We accomplish this objective through a bidirectional approach: Firstly, delaying the onset of the roofline turning point, and secondly, concurrently reducing the computation intensity within the fused kernel. For the first part, we replace all FP32 operations in the original TensorRT-LLM kernel with their FP16 counterpart, effectively doubling the computation roof. For the second part, we observe that the arithmetic intensity of dequantization can be significantly reduced to 2 operations per element by applying bit tricks proposed in [20]. Furthermore, we note that simplifying the control logic and prefetching the scaling factors and zero values, thereby simplifying address calculations, contribute to performance improvements. After incorporating these enhancements, we observe a 1.5$\times$ speedup over TensorRT-LLM’s KV8 kernel on A100.
VI Evaluation
VI-A Evaluation Setup
Algorithm
The QoQ quantization algorithm is implemented using HuggingFace [37] on top of PyTorch [26]. We use per-channel symmetric INT8 quantization on activations, and per-token asymmetric INT4 group quantization on KV cache. “W4A8KV4 g128” refers to the case where QServe used progressive group quantization on weights: per-channel symmetric INT8 quantization followed by asymmetric INT4 quantization with a group size of 128, while “W4A8KV4” is the per-channel counterpart for weight quantization.
System
QServe serving system is implemented using CUDA and PTX assembly for high-performance GPU kernels. We also provide a purely PyTorch-based front-end framework for better flexibility. For the throughput benchmarking, we perform all experiments under PyTorch 2.2.0 with CUDA 12.2, unless otherwise specified. The throughput numbers reported are real measurements on NVIDIA GPUs. For baseline systems, we use TensorRT-LLM v0.9.0 and latest main branches from QuaRot and Atom as of April 18th, 2024. Paged attention is enabled for all systems except QuaRot, which does not offer corresponding support.
VI-B Accuracy Evaluation
Benchmarks
We evaluated QoQ on the Llama-1 [33], Llama-2 [34], Llama-3 families, Mistral-7B [17], Mixtral-8x7B [18] and Yi-34B [39] models. Following previous literature [8, 12, 44, 2, 38, 23], we evaluated QoQ-quantized models on language modeling and zero-shot tasks. Specifically, we evaluated on WikiText2 [24] for perplexity, and evaluated on PIQA [3] (PQ), ARC [5], HellaSwag [42] (HS) and WinoGrande [29] (WG) with lm_eval [13].
Baselines
We compared QoQ to widely used post-training LLM quantization techiniques, SmoothQuant [38], GPTQ [12], AWQ [23], and recently released state-of-the-art 4-bit weight-activation quantization frameworks, Atom [44] and QuaRot [2]. For SmoothQuant, we uses static per-tensor symmetric 8-bit quantization for KV cache following the settings in the TensorRT-LLM [25]. For GPTQ, we use their latest version with “reorder” trick, denoted as “GPTQ-R”. For QuaRot and Atom, we mainly evaluated using Pile validation dataset as calibration dataset. We also report their results with WikiText2 as calibration dataset in gray color. For “W4A8KV4 g128” setting, both QuaRot and Atom does not support progressive group quantization, and thus we evaluated them using ordinary group weight quantization (i.e., each group has one FP16 scale factor). Unsupported models and quantization settings will be reported as NaN.
@@ VI-B Accuracy Evaluation
TABLE II: WikiText2 perplexity with 2048 sequence length. The lower is the better.
@@
$\ddagger$ QuaRot and Atom use ordinary group quantization where each group has one FP16 scale factor.
TABLE III: Zero-shot accuracy on five common sense tasks with 2048 sequence length.
@@
accuracy from 72.9 to 63.7 on ARC easy and from 44.7 to 35.5 on ARC challenge task. [28]
WikiText2 perplexity
Table II compares the Wikitext2 perplexity results between QoQ and other baselines. For Llama-2-7B, compared to W8A8 SmoothQuant and W4A16 AWQ, QoQ only increased perplexity by up to 0.16 QoQ consistently outperformed Atom with either W4A4 or W4A8KV4 quantization precision. QoQ also showed up to 0.49 perplexity improvement compared to W4A4 Quarot.
Zero-shot accuracy
we report the zero-shot accuracy of five common sense tasks in Table III. QoQ significantly outperformed other 4-bit quantization methods. Especially on the Winogrande task, compared to Quarot, QoQ accuracy is 4.82% higher. Compared to FP16, QoQ only introduced 1.03%, 0.89% and 0.40% accuracy loss for Llama-2 at 7B, 13B and 70B size.
Figure 15: QServe significantly outperforms existing large language model (LLM) serving frameworks in batched generation tasks across different LLMs, ranging from 7B to 72B models. It achieves an average speedup of 2.36$\times$ over the state-of-the-art LLM serving system, TensorRT-LLM v0.9.0, on the L40S GPU, and it is also 1.68$\times$ faster on the A100 GPU. All experiments were conducted under the same device memory budget (i.e. 80GB on A100 and 48GB on L40S). We omit the geometric mean speedup of Atom since it only supports Llama2-7B. For absolute values, see Table IV.
TABLE IV: The absolute token generation throughput of QServe and TensorRT-LLM in Fig. 15. *: we calculate the speedup over highest achieveable throughput from TensorRT-LLM across all three precision configurations. Our QServe system achieves competitive throughput on L40S GPU compared to TensorRT-LLM on A100, effectively reducing the dollar cost of LLM serving by 3$\times$. Unit: tokens/second.
@@
VI-C Efficiency Evaluation
We present relative performance comparisons in Figure 15 and absolute throughput values in Table IV. We use per-channel quantization for A100 and per-group quantization for L40S. This is because L40S has stronger CUDA cores for dequantization. Relative to the best-performing configuration of TensorRT-LLM, QServe demonstrates significant improvements on A100: it achieves 2$\times$ higher throughput for Llama-1-30B, 1.2-1.4$\times$ higher throughput for Llama-2 models, 1.2$\times$ higher throughput for Mistral and Yi, and 2.4$\times$ higher throughput for Qwen-1.5. The performance improvements are particularly notable on the L40S GPUs, where we observed a throughput improvement ranging from 1.47$\times$ to 3.47$\times$ across all seven models evaluated. Remarkably, despite the L40S’s significantly smaller memory capacity compared to the A100, QServe effectively maintains the same batch size as TensorRT-LLM on the A100. This achievement is attributed to our aggressive 4-bit quantization applied to both weights and the KV cache. By examining Table IV, we clearly observe that serving five of seven models under 34B on L40S with QServe achieves even higher throughput than serving them on A100 using TensorRT-LLM. Our performance gain over Atom and QuaRot on A100 is even more prominent since these systems did not outperform TensorRT-LLM. On L40S, QServe still achieves 10% higher throughput than Atom when running Llama-2-7B, the only model supported by their system despite the fact that we use higher quantization precision. Besides, the accuracy achieved by QServe is much better than Atom, as indicated in Table III.
VI-D Analysis and Discussion.

### v3

@@ 4.2 SmoothAttention
As illustrated in Figure 16, directly reducing the KV cache to 4 bits significantly degrades the LLM accuracy. We visualize the magnitude distributions of the sampled Key and Value cache activations in Figure 7. We observe that: the Value matrices show no significant outlier pattern, whereas Key matrices tend to have fixed outlier channels in each head. These outliers are $\sim$10$\times$ larger than most of activation values. Though they can be easily handled KV8 quantization in prior works Xiao et al. (2023), it places challenging obstacle to KV4 quantization due to less quantization levels.
Inspired by SmoothQuant Xiao et al. (2023), we propose SmoothAttention to scale down the outlier channels in Key cache by a per-channel factor $\mathbf{\lambda}$:
$\mathbf{Z}=\left(\mathbf{Q}\mathbf{\Lambda}\right)\cdot\left(\mathbf{K}\mathbf{\Lambda}^{-1}\right)^{T},\;\;\;\mathbf{\Lambda}=\mathrm{diag}\left(\mathbf{\lambda}\right)$ (7)
@@
$\mathbf{\lambda}_{i}=\max\left(|\mathbf{K}_{i}|\right)^{\alpha}.$ (8)
@@
$\mathbf{\lambda}_{i}=\lambda_{i+\frac{D}{2}}=\max\left(\max\left(|\mathbf{K}_{i}|\right),\max\left(|\mathbf{K}_{i+\frac{D}{2}}|\right)\right)^{\alpha}$ (9)
@@
4.3 General LLM Quantization Optimizations
@@
4.3.1 Block Input Module Rotation
In transformer blocks, we define the components that take in the block inputs as input modules, such as the QKV Projection Layer and the FFN 1st Layer. As shown in Figure 9, inspired by Ashkboos et al. (2024); Chee et al. (2024), we rotate the block input activations by multiplying the rotation matrix. To keep mathematical equivalence of linear layers, we rotate the corresponding weights accordingly in the reversed direction. After rotation, each channel’s activations are linear combinations of all other channels, and thus outlier channels are effectively suppressed. Furthermore, since rotation is a unitary transformation, we can fuse the rotation matrix with the previous linear layers’ weights. We simply choose the scaled Hadamard matrix as the rotation matrix.
4.3.2 Block Output Module Smoothing
Output modules refer to those layers that generate block outputs, such as the Output Projection Layer and FFN 2nd Layer. As shown in Figure 9, inspired by Xiao et al. (2023), we smooth the block intermediate activations through dividing them by a per-channel smoothing factor. Original SmoothQuant does not smooth the block intermediate activations; moreover, if we directly smooth these modules with the same migration strength as input modules (e.g., q_proj, up_proj), the evaluated Wikitext-2 perplexity of the Llama-2-7B model will drop by as much as 0.05. In practice, we find that the migration strength $\alpha$ should be near 0. That is, the smoothing factor $\lambda$ is mostly determined by weights instead of activations, which is very different from the observations in SmoothQuant.
4.3.3 Activation-Aware Channel Reordering
@@
Both AWQ Lin et al. (2024) and Atom Zhao et al. (2023) have observed that maintaining the salient weights in FP16 can significantly improve model accuracy. These salient weights can be identified by the activation distribution. Instead of introducing mixed-precision quantization used by Atom, we propose activation-aware channel reordering as shown in Figure 10. We use $\max\left(|\mathbf{X}|\right)$ to determine the channel salience, and then reorder channels so that channels with similar salience are in the same quantization group.
4.3.4 Weight Clipping
Weight clipping is another popular quantization optimization technique. It applies a clip ratio $\alpha$ to the dynamic range in Equation 2 by letting $\mathbf{W}_{\max}=\alpha\max\left(\mathbf{W}\right)$ and $\mathbf{W}_{\min}=\alpha\min\left(\mathbf{W}\right)$. Previous approaches Frantar et al. (2022); Zhao et al. (2023); Ashkboos et al. (2024); Lin et al. (2024) grid search the clip ratio $\alpha$ to minimize either quantization error of tensor itself (i.e., $\|\mathbf{W}-Q\left(\mathbf{W};\alpha\right)\|$) or output mean square error (i.e., $\|\mathbf{X}\mathbf{W}^{T}-\mathbf{X}Q\left(\mathbf{W}^{T};\alpha\right)\|$. In QoQ, we minimize the layer output error for all linear layers, expect for q_proj and k_proj, for which we optimize block output mean square error:
$\arg\min_{\alpha}\|\mathrm{Block}\left(\mathbf{X};\mathbf{W}\right)-\mathrm{Block}\left(\mathbf{X};Q\left(\mathbf{W};\alpha\right)\right)\|.$ (10)
5 QServe Serving System
To this end, we have presented the QoQ quantization algorithm, which aims to minimize accuracy loss incurred by W4A8KV4 quantization. However, realizing the theoretical throughput benefits in Figure 3 remains challenging. Thus, in this section, we will delve into the QServe system design, which follows two important directions: I. Reducing main loop overhead in GEMM kernels; II. Accerating KV4 attention.
5.1 QServe System Runtime
@@
KV Cache Management. To avoid memory fragmentation, we follow vLLM Kwon et al. (2023a) and TensorRT-LLM NVIDIA (2023) to adopt paged KV caches. In contrast to these two frameworks, which perform per-tensor, static quantization (i.e., scaling factors computed offline) on KV caches, QServe requires per-head, dynamic KV quantization to maintain competitive accuracy due to the lower bit precision (4 vs. 8). We therefore store FP16 scaling factors and zero points for each head immediately following the quantized KV features in each KV cache page, allowing these values to be updated on-the-fly. QServe also supports in-flight batching, similar to vLLM and TensorRT-LLM.
5.2 W4A8 GEMM in QServe
As discussed in Section 3, the main loop overhead poses a significant obstacle in allowing quantized GEMMs to attain the theoretical performance gains projected by the roofline model (Figure 3). Therefore, the focus of QServe W4A8 GEMM is to reduce main loop overhead. Specifically, we address the costs of pointer arithmetic operations through compute-aware weight reorder, and reduce dequantization overhead through a subtraction after multiplication computation order and register-level parallelism.
5.2.1 Compute-Aware Weight Reorder
@@
5.2.2 Fast Dequantization in Per-Channel W4A8 GEMM
@@
$\mathbf{O}=\hat{\mathbf{X}}\hat{\mathbf{W}}=(\mathbf{Q}_{\mathbf{X}}\odot\mathbf{S}_{\mathbf{X}})((\mathbf{Q}_{\mathbf{W}}-\mathbf{Z}_{\mathbf{W}})\odot\mathbf{S}_{\mathbf{W}}),$ (11)
@@
$\mathbf{O}$ $=(\mathbf{Q}_{\mathbf{X}}\odot\mathbf{S}_{\mathbf{X}})(\mathbf{Q}_{\mathbf{W}}\odot\mathbf{S}_{\mathbf{W}}-\mathbf{ZS}_{\mathbf{W}})$ (12)
@@
5.2.3 Fast Dequantization in Per-Group W4A8 GEMM
@@ 5.2.4 General Optimizations
5.3 KV4 Attention in QServe
Table 1: A naive KV4 attention implementation is 1.7$\times$ faster on L40S than TRT-LLM-KV8, but is 1.1-1.2$\times$ slower on A100 due to earlier CUDA core roofline turning point.
@@ 5.3 KV4 Attention in QServe
We start with the KV8-attention decoding stage kernel from TensorRT-LLM as our baseline and replace all static, per-tensor quantized 8-bit KV cache accesses and conversions with their dynamic, per-head quantized 4-bit counterparts. This direct replacement immediately leads to 1.7$\times$ speedup on L40S, but results in 1.2$\times$ slowdown on A100 (Table 1), compared to the KV8 baseline.
@@
Once again, our analysis reveals that the devil is in the slow CUDA cores, which are responsible for executing the attention kernels during the decoding stage. While each individual batched GEMV has a computation intensity of 1 MAC / element, the computation intensity escalates significantly for a fused attention kernel that combines all the arithmetics and KV cache updates. As an illustration, naively dequantizing a single INT4 number from the KV cache necessitates 5 ALU Ops. This includes mask and shift operations to isolate the operand, type conversion from integer to floating-point representation, and floating point mul and sub to obtain the final results. It is crucial to note that the roofline turning point for A100 FP32 CUDA cores is merely 9.8 Ops/Byte. That said, the dequantization of KV operands alone already saturates this bound, leading to the surprising observation that the fused KV4 attention kernel can become compute-bound on datacenter GPUs like A100. In fact, similar observations hold in other systems like QuaRot Ashkboos et al. (2024) and Atom Zhao et al. (2023). Specifically, QuaRot introduces compute-intensive Hadamard transformation Chee et al. (2024) in the attention operator, making it hard to achieve real speedup over TRT-LLM-KV8 with 4-bit quantized KV caches.
@@
To mitigate the compute-bound bottleneck, it is important to shift the decoding stage KV4 attention kernels away from the compute-bound region. We accomplish this objective through a bidirectional approach: Firstly, delaying the onset of the roofline turning point, and secondly, concurrently reducing the computation intensity within the fused kernel. For the first part, we replace all FP32 operations in the original TensorRT-LLM kernel with their FP16 counterpart, effectively doubling the computation roof. For the second part, we observe that the arithmetic intensity of dequantization can be significantly reduced to 2 operations per element by applying bit tricks proposed in Kim et al. (2022). Furthermore, we note that simplifying the control logic and prefetching the scaling factors and zero values, thereby simplifying address calculations, contribute to performance improvements. After incorporating these enhancements, we observe a 1.5$\times$ speedup over TensorRT-LLM’s KV8 kernel on A100.
@@
Table 2: WikiText2 perplexity with 2048 sequence length. The lower is the better.
@@
Table 3: Zero-shot accuracy on five common sense tasks with 2048 sequence length.
@@
accuracy from 72.9 to 63.7 on ARC easy and from 44.7 to 35.5 on ARC challenge task. Rouhani et al. (2023)
Figure 15: QServe significantly outperforms existing large language model (LLM) serving frameworks in batched generation tasks across different LLMs, ranging from 7B to 72B models. It achieves an average speedup of 2.36$\times$ over the state-of-the-art LLM serving system, TensorRT-LLM v0.9.0, on the L40S GPU, and it is also 1.68$\times$ faster on the A100 GPU. All experiments were conducted under the same device memory budget (i.e. 80GB on A100 and 48GB on L40S). We omit the geometric mean speedup of Atom since it only supports Llama2-7B. For absolute values, see Table 4.
Table 4: The absolute token generation throughput of QServe and TensorRT-LLM in Fig. 15. *: we calculate the speedup over highest achieveable throughput from TensorRT-LLM across all three precision configurations. Our QServe system achieves competitive throughput on L40S GPU compared to TensorRT-LLM on A100, effectively reducing the dollar cost of LLM serving by 3$\times$. Unit: tokens/second.
@@
6 Evaluation
6.1 Evaluation Setup
Algorithm. The QoQ quantization algorithm is implemented using HuggingFace Wolf et al. (2020) on top of PyTorch Paszke et al. (2019). We use per-token symmetric INT8 quantization on activations, and per-head asymmetric INT4 quantization on KV cache. “W4A8KV4 g128” refers to the case where QServe used progressive group quantization on weights: per-channel symmetric INT8 quantization followed by asymmetric INT4 quantization with a group size of 128, while “W4A8KV4” is the per-channel counterpart for weight quantization.
System. QServe serving system is implemented using CUDA and PTX assembly for high-performance GPU kernels. We also provide a purely PyTorch-based front-end framework for better flexibility. For the throughput benchmarking, we perform all experiments under PyTorch 2.2.0 with CUDA 12.2, unless otherwise specified. The throughput numbers reported are real measurements on NVIDIA GPUs. For baseline systems, we use TensorRT-LLM v0.9.0 and latest main branches from QuaRot and Atom as of April 18th, 2024. Paged attention is enabled for all systems except QuaRot, which does not offer corresponding support.
6.2 Accuracy Evaluation
Benchmarks.
We evaluated QoQ on the Llama-1 Touvron et al. (2023a), Llama-2 Touvron et al. (2023b), Llama-3 families, Mistral-7B Jiang et al. (2023), Mixtral-8x7B Jiang et al. (2024) and Yi-34B Young et al. (2024) models. Following previous literature Dettmers et al. (2022); Frantar et al. (2022); Zhao et al. (2023); Ashkboos et al. (2024); Xiao et al. (2023); Lin et al. (2024), we evaluated QoQ-quantized models on language modeling and zero-shot tasks. Specifically, we evaluated on WikiText2 Merity et al. (2016) for perplexity, and evaluated on PIQA Bisk et al. (2020) (PQ), ARC Clark et al. (2018), HellaSwag Zellers et al. (2019) (HS) and WinoGrande Sakaguchi et al. (2019) (WG) with lm_eval Gao et al. (2023).
Baselines.
We compared QoQ to widely used post-training LLM quantization techiniques, SmoothQuant Xiao et al. (2023), GPTQ Frantar et al. (2022), AWQ Lin et al. (2024), and recently released state-of-the-art 4-bit weight-activation quantization frameworks, Atom Zhao et al. (2023) and QuaRot Ashkboos et al. (2024). For SmoothQuant, we uses static per-tensor symmetric 8-bit quantization for KV cache following the settings in the TensorRT-LLM NVIDIA (2023). For GPTQ, we use their latest version with “reorder” trick, denoted as “GPTQ-R”. For QuaRot and Atom, we mainly evaluated using Pile validation dataset as calibration dataset. We also report their results with WikiText2 as calibration dataset in gray color. For “W4A8KV4 g128” setting, both QuaRot and Atom does not support progressive group quantization, and thus we evaluated them using ordinary group weight quantization (i.e., each group has one FP16 scale factor). Unsupported models and quantization settings will be reported as NaN.
WikiText2 perplexity.
Table 2 compares the Wikitext2 perplexity results between QoQ and other baselines. For Llama-2-7B, compared to W8A8 SmoothQuant and W4A16 AWQ, QoQ only increased perplexity by up to 0.16 QoQ consistently outperformed Atom with either W4A4 or W4A8KV4 quantization precision. QoQ also showed up to 0.49 perplexity improvement compared to W4A4 Quarot.
Zero-shot Accuracy and Long-Context Accuracy.
We report the zero-shot accuracy of five common sense tasks in Table 3. QoQ significantly outperformed other 4-bit quantization methods. Especially on the Winogrande task, compared to Quarot, QoQ accuracy is 4.82% higher. Compared to FP16, QoQ only introduced 1.03%, 0.89% and 0.40% accuracy loss for Llama-2 at 7B, 13B and 70B size. Furthermore, our results in Table 5 demonstrate that QoQ can maintain minimal degradation on the long-context performance relative to the BF16 baseline.
Table 5: LongBench evaluation for Llama-3.1-8b-Instruct using QoQ W4A8KV4 g128.
DuReader
GovReport
HotpotQA
MultiNews
Musique
QMSum
SamSum
TriviaQA
TREC
MultiFieldQA-En
Average
BF16 35.07 34.54 16.68 26.84 11.68 23.48 43.50 91.65 72.50 29.22 38.52
QoQ 35.45 34.09 17.46 26.73 12.05 23.45 44.42 91.45 71.00 27.65 38.38